OpenStudy (anonymous):

An equation and the steps to solve it are shown below:

OpenStudy (quin100):

were is the equation

OpenStudy (anonymous):

x=0? lol ;) jk awaiting edit

OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 3x - one over two (8x - 2) = 4 Step 1: 3x - 4x + 1 = 4 Step 2: x + 1 = 4 Step 3: x + 1 - 1 = 4 - 1 Step 4: x = 3 Which step was the error made and why? Step 1, because incorrect signs have been used Step 2, because the terms have been added incorrectly Step 3, because numbers have been added to both sides of the equation No error was made

OpenStudy (anonymous):

combine 3x and -4x and then continue...

OpenStudy (anonymous):

so you will have -1x + 1 = 4.. then solve...

OpenStudy (anonymous):

x=-3